OpenStudy (anonymous):

What is an equation of a line that is parallel to y = 4x – 1 and passes through (2, 15)?

OpenStudy (mathstudent55):

A line that is parallel to another hase the same slope and different y-intercept. y = 4x - 1 has slope = 4 and y-itercept = -1, or point (0,-1) The parallel line will have the same slope, so it will be y = 4x + b. Now you need to find b. Since you have a point on that line, (2,15), insert (2,15) in for x and y of y = 4x + b and solve for b. Once you know b, write it as y= 4x + b ( replace b with its value)

OpenStudy (anonymous):

So the answer is y=4x + ? Ugh. How am I supposed to solve for b?

OpenStudy (mathstudent55):

Read second part of my answer above, from 4th line on.

OpenStudy (mathstudent55):

Insert (2,15) into y = 4x + b. That is, solve y = 4x + b with x = 2 and y = 15, and solve for b.

OpenStudy (anonymous):

Ok so 15= 4x12 ?

OpenStudy (mathstudent55):

y = 4x + b 15 = 4(2) + b 15 = 8 + b b + 8 = 15 b = 7 y = 4x + 7 final answer

OpenStudy (anonymous):

So B=7?

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

Thank you so much

OpenStudy (mathstudent55):

wlcm